OpenStudy (anonymous):

if the permeter of a rectangle is 60 meters and its area is 200 square meter then one of it's sides must be ------ the other? options: A=one third of B=triple C=double D=one tenth of

Parth (parthkohli):

Do you know how to solve simultaneous equations?

OpenStudy (anonymous):

yes

Parth (parthkohli):

So the equation says that if 'a' and 'b' are the two adjacent sides, then: ab = 200 2(a + b) = 60

OpenStudy (anonymous):

draw out a rectangle you should have 2 lengths and 2 widths just remember perimeter = l + l + w + w and area = length times width

Parth (parthkohli):

@laraib Can you solve the equation I gave ya?

OpenStudy (anonymous):

yes m solving

Parth (parthkohli):

If I were you, I'd be using the eyeball method.

OpenStudy (anonymous):

how? tell me plz

Parth (parthkohli):

Two numbers that add up to 30 and multiply to get 200.

Parth (parthkohli):

Clearly, those numbers are 20 and 10 right?

OpenStudy (anonymous):

yes that's how i solved it

Parth (parthkohli):

very good! So now look at the choices you have in your question.

OpenStudy (anonymous):

it would be DOUBLE.right?

Parth (parthkohli):

Yes :)

OpenStudy (anonymous):

thanks. can u plz suggest me any website from where i can prepare for aptitude test of universities?

Parth (parthkohli):

http://openstudy.com :)

OpenStudy (anonymous):

but here ill get the solution of those questions which i have i want to prepare topic wise

Parth (parthkohli):

You can find a good book.

OpenStudy (anonymous):

here on open study?

Parth (parthkohli):

Nope, not here... search for an ebook.

OpenStudy (anonymous):

if i^2= -1 then i^23+i^12+i^11 /10 =?

OpenStudy (anonymous):

what is ebook?

Parth (parthkohli):

This?\[\rm {i^{23} + i^{12}+{i^{11} }\over 10}\]

OpenStudy (anonymous):

yup

Parth (parthkohli):

Okay, factor.

OpenStudy (anonymous):

it's 11 i guess not 23

OpenStudy (anonymous):

there is a misprint actually

Parth (parthkohli):

?

OpenStudy (anonymous):

i^11 instead of i^23

Parth (parthkohli):

So this?\[\rm {i^{11} + i^{12}+i^{11}\over 10}\]

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yup :)

Parth (parthkohli):

Can you factor it?

OpenStudy (anonymous):

yes

Parth (parthkohli):

So do it. :P

OpenStudy (anonymous):

but i didnt get the correct answer

Parth (parthkohli):

Can you just show me how you factored it?

OpenStudy (anonymous):

you give me the solution

Parth (parthkohli):

?

OpenStudy (anonymous):

i got -2i+1 / 10

Parth (parthkohli):

\[\rm {i^{11}(1 + 1 + i)}\]

Parth (parthkohli):

Can you figure out what \(\rm i^{11}\) is?

OpenStudy (anonymous):

options are 0.1 0 -1 1

OpenStudy (anonymous):

yup it's -i

Parth (parthkohli):

Yes.

Parth (parthkohli):

You have\[\rm {-i(2 + i) \over 10}\]

Parth (parthkohli):

None of the choices...

hartnn (hartnn):

if only real part is asked then there is an option

Parth (parthkohli):

Yep.

OpenStudy (anonymous):

ok

Parth (parthkohli):

@laraib Use distributive property.

OpenStudy (anonymous):

if two vertices of an equilateral triangle are (2,3) and (5,3) then the area of the triangle is?